the enthalpy of formation of Methane carbon dioxide and water are - 74.8 393.5 , - 286.2 kJ respectively calculate the enthalpy of combustion of Methane at ordinary temperature?

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 

CH4 + 2O2 → CO2 +2 H2O this is the reaction of combustion of methane.

C + O2 → CO2 ∆H1 = -393 Kj/mol
H2 + 1/2O2→ H2O; ∆ H2 = 286.2 Kj/mol 
C + 2H2 → CH4 ; ∆H3 = -74.8 kj/mol 

CH4 → C + 2H2 , ∆H3' = 74.8 
C + O2 → CO2 ∆ H1 = -393 
2( H2 + 1/2O2 → H2O ; ∆H2 = 2× 286.2 

=============================
add this reactions, 

CH4 + 2O2 → CO2 +2H2O 

∆H =( 74.6 -393 +2×286.2 ) Kj/mol 
=254 Kj/mol 

 
Regards
Arun (askIItians forum expert)